### v2

@@ Abstract
keywords:
cite=DBLP:conf/stacs/Main24
@@ 1 Introduction
In some applications, this purely adversarial model may be too restrictive. This is the case in settings with several agents, each with their own objective, who are not necessarily opposed to one another. Such situations are modelled by multiplayer non-zero-sum games on graphs. The counterpart of optimal strategies in this setting is typically a notion of equilibrium. We focus on Nash equilibria [29] (NEs) in the following; an NE is a tuple of strategies, one per player, such that no player has an incentive to unilaterally deviate from their strategy.
@@
We study games on possibly infinite arenas in which all players share objectives of the same type. We focus on variants of reachability objectives; reachability objectives constitute a fundamental class of objectives in reactive synthesis [12].
A reachability objective models the goal of visiting a given set of target vertices. The complement of a reachability objective is called a safety objective: it requires that a set of bad vertices be avoided. A Büchi objective requires that a target set be visited infinitely often, and the complement of a Büchi objective is called a co-Büchi objective. We also study games with a quantitative variant of the reachability objective. We consider arenas with edges labelled by non-negative integer weights (modelling, e.g., the passage of time), and define the cost of a player as the sum of weights prior to the first visit of a target, or infinity if the target is never reached. This cost function is called the shortest-path cost function. The goal of a player is to minimise it (i.e., to reach a target as soon as possible).
NEs are guaranteed to exist in games with these objectives or cost functions: e.g., see [10, 17] for reachability games and for shortest-path games in finite arenas, B for shortest-path games in general and [32] for safety, Büchi and co-Büchi games.
Usually, finite-memory NEs for these games are given by strategies whose size depends on the arena (e.g., [10, 33, 9]). These constructions consequently do not generalise to infinite arenas. The main idea of these approaches is as follows. First, one shows that there exist plays resulting from NEs with a finite representation, e.g., a lasso. This play is then encoded in a Mealy machine. If some player is inconsistent with the play, the other players switch to a (finite-memory) punishing strategy to sabotage the deviating player; this enforces the stability of the equilibrium. This punishing mechanism is inspired by the proof of the folk theorem for NEs in repeated games [20, 30].
@@
Second, for reachability, shortest-path and safety games, we provide memory bounds that are independent of the size of the arena and depend only on the number of players. For Büchi and co-Büchi games, we show that similar arena-independent memory bounds cannot be obtained: we provide a family of two-player games where NEs with an outcome in which the second player wins require a memory of size linear in the size of the arena (Example 7.3). We summarise our memory upper bounds for each considered type of game in Table 1.
@@
Reachability $n^{2}$ (Thm. 5.10)
Shortest-path $n^{2}+2n$ (Thm. 5.13)
@@
Büchi $|V|+n^{2}+n$ (Thm. 7.6) Finite-memory (Thm. 7.67.11)
Co-Büchi $|V|+2n$ (Thm. 7.15) Finite-memory (Thm. 7.15)
@@
Our key observation is that it is not necessary to fully implement the punishment mechanism: some deviations do not warrant switching to punishing strategies. This allows us to encode only part of the information in the memory instead of an entire play.
@@
We refer to the survey [13] for an extensive bibliography on games played on finite graphs, to [12] for a survey centred around reachability games and to [19] as a general reference on games on graphs. We discuss three research direction related to this work.
@@ 2 Preliminaries
Let $(V,E)$ be a directed graph where $V$ is a (possibly infinite) set of vertices and $E\subseteq V\times V$ is an edge relation. For any $v\in V$, we write $\mathsf{Succ}_{E}(v)=\{v^{\prime}\in V\mid(v,v^{\prime})\in E\}$ for the set of successor vertices of $v$. Let $n\in\mathbb{N}$. An $n$-player arena (on $(V,E)$) is a tuple $\mathcal{A}=((V_{i})_{i\in\llbracket{}n\rrbracket{}},E)$, where $(V_{i})_{i\in\llbracket{}n\rrbracket{}}$ is a partition of $V$. We assume that there are no deadlocks in the arenas we consider, i.e., for all $v\in V$, $\mathsf{Succ}_{E}(v)$ is not empty. We write $\mathcal{P}_{i}$ for player $i$.
@@
A play or history is simple if no vertex occurs twice (or more) within. A cycle is a history of the form $h=h^{\prime}\mathsf{first}(h^{\prime})$ where $h^{\prime}$ is a history, and $h=h^{\prime}\mathsf{first}(h^{\prime})$ is a simple cycle if $h^{\prime}$ is a simple history. A play is a lasso if it is of the form $pc^{\omega}$ where $pc\in\mathsf{Hist}(\mathcal{A})$, and we say that a lasso $pc^{\omega}$ is simple if $pc$ is a simple history
A segment of a play $\pi$ is either a suffix $\pi{\ell}$ of $\pi$ ($\ell\in\mathbb{N}$) or any history of the form $v_{\ell}\ldots v_{\ell^{\prime}}$ where $\ell\leq\ell^{\prime}$ (i.e., an infix of $\pi$). We denote segments by $\mathsf{sg}$ to avoid distinguishing finite and infinite segments of plays in the following. A segment is simple if it is a simple history, a simple play or a simple lasso. A segment is trivial if it consists of a single vertex.
@@
We identify two classes of strategies of interest in this work. A strategy $\sigma_{i}$ is memoryless if the moves it prescribes depend only on the current vertex, i.e., if for all $h$, $h^{\prime}\in\mathsf{Hist}_{i}(\mathcal{A})$, if $\mathsf{last}(h)=\mathsf{last}(h^{\prime})$, then $\sigma_{i}(h)=\sigma_{i}(h^{\prime})$. We view memoryless strategies as functions $V_{i}\to V$.
@@
To describe the strategy induced by a Mealy machine, we first define the iterated update function $\widehat{\mathsf{up}}\colon V^{*}\to M$ by induction. We write $\varepsilon$ for the empty word. We let $\widehat{\mathsf{up}}(\varepsilon)=m_{\mathsf{init}}$ and for all $wv\in V^{*}$, $\widehat{\mathsf{up}}(wv)=\mathsf{up}(\widehat{\mathsf{up}}(w),v)$. The strategy $\sigma_{i}^{\mathcal{M}}$ induced by $\mathcal{M}$ is defined, for all histories $h=h^{\prime}v\in\mathsf{Hist}_{i}(\mathcal{A})$, by $\sigma_{i}^{\mathcal{M}}(h)=\mathsf{nxt}_{i}(\widehat{\mathsf{up}}(h^{\prime}),v)$.
We say that a finite-memory strategy $\sigma_{i}$ has memory size at most $b\in\mathbb{N}$ if there is some Mealy machine $(M,m_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$ encoding $\sigma_{i}$ with $|M|\leq b$.
@@
In finite arenas, these two variants of Mealy machines share the same expressiveness. Given an edge-update Mealy machine, we can transform it into an equivalent vertex-update Mealy machine by augmenting the memory states with the last vertex to have been visited. With this last vertex and the current one, the last edge can be inferred. This allows memory updates to be imitated one step later than in the original edge-update Mealy machine. In general, edge-update Mealy machines may be more concise than vertex-update ones as they have access to more information in memory updates. It can be shown that transforming an edge-update Mealy machine into an equivalent vertex-update strategy may induce an increase in size by a multiple of the arena size in the worst case.
@@
We formalise the goal of a player in two ways. In the qualitative case, we describe the goal of a player by a set of plays, called an objective. We say that a play $\pi$ satisfies an objective $\Omega$ if $\pi\in\Omega$. For quantitative specifications, we assign a quantity to each play using a cost function $\mathsf{cost}_{i}\colon\mathsf{Plays}(\mathcal{A})\to\overline{\mathbb{R}}$ that $\mathcal{P}_{i}$ intends to minimise. Any goal expressed by an objective $\Omega$ can be encoded using a cost function $\mathsf{cost}_{i}$ which assigns 00 to plays in $\Omega$ and $1$ to others (i.e., the indicator of $\mathsf{Plays}(\mathcal{A})\setminus\Omega$); aiming to minimise this cost is equivalent to aiming to satisfy the objective. For this reason, we present further definitions using cost functions, and explicitly mention when notions are specific to objectives.
@@
We study five classes of games with goals related to reachability. We consider a qualitative and quantitative formulation for the goal of reaching a target, and qualitative formulations of the goals of avoiding a target, of infinitely often reaching a target and of avoiding a target from some point on. Let $T\subseteq V$ denote a set of target vertices. We refer to the set of target vertices $T$ as a target.
We first define the reachability objective, which models the goal of reaching $T$. Formally, the reachability objective (for $T$) $\mathsf{Reach}(T)$ is defined by the set $\{v_{0}v_{1}v_{2}\ldots\in\mathsf{Plays}(\mathcal{A})\mid\exists\,\ell\in\mathbb{N},\,v_{\ell}\in T\}$. The complement of the reachability objective $\mathsf{Safe}(T)=\mathsf{Plays}(\mathcal{A})\setminus\mathsf{Reach}(T)$, which formalises the goal of avoiding $T$, is called the safety objective (for $T$).
Next, we introduce a cost function formalising the goal of reaching a target as soon as possible. In this context, we assign (non-negative integer) weights to edges via a weight function $w\colon E\to\mathbb{N}$, which model, e.g., the time taken when traversing an edge. The weight function is extended to histories as follows: for $h=v_{0}\ldots v_{r}\in\mathsf{Hist}(\mathcal{A})$, we let $w(h)=\sum_{\ell=0}^{r-1}w((v_{\ell},v_{\ell+1}))$. We define the shortest-path cost function (for $T$ and $w$), for all plays $\pi=v_{0}v_{1}\ldots$, by $\mathsf{SPath}^{T}_{w}(\pi)=w(\pi{r})$ if $r=\min\{\ell\in\mathbb{N}\mid v_{\ell}\in T\}$ exists and $\mathsf{SPath}^{T}_{w}(\pi)=+\infty$ otherwise. By assigning an infinite cost to plays that do not visit $T$, we model the idea that not reaching $T$ is the least desirable situation.
Finally, we define the Büchi objective, expressing the goal of reaching a target infinitely often. Formally, the Büchi objective (for $T$) $\mathsf{B\ddot{u}chi}(T)$ is defined by the set $\{v_{0}v_{1}v_{2}\ldots\in\mathsf{Plays}(\mathcal{A})\mid\forall\,\ell\in\mathbb{N},\,\exists\,\ell^{\prime}\geq\ell,\,v_{\ell^{\prime}}\in T\}$. The complement of a Büchi objective is a co-Büchi objective: the co-Büchi objective (for $T$), which expresses the goal of visiting $T$ finitely often, is defined as $\mathsf{coB\ddot{u}chi}(T)=\mathsf{Plays}(\mathcal{A})\setminus\mathsf{B\ddot{u}chi}(T)$.
@@
NEs are guaranteed to exist in the five classes of games we consider; see, e.g. [10, 17] for reachability games and for shortest-path games in finite arenas, B for shortest-path games in general and [32] for safety, Büchi and co-Büchi games. Even though NEs exist from all initial vertices, there are no guarantees on the quality of their cost profiles (this is a core motivation of the constrained NE existence problem mentioned in Section 1). In a given game, there can exist several NEs from a given initial vertex with incomparable cost profiles. Similarly, in a given game with objectives, from a fixed initial vertex, there can exist NEs with which all players satisfy their objective and NEs with which no players satisfy their objective [32, 33]. In general, memory may be necessary to construct an NE with a cost profile that is preferable to a given vector (i.e., a solution to the constrained existence problem).
@@
[…]
(a) A two-player weighted arena in which there two memoryless NE with incomparable payoff profiles. Edges are labelled by their weight and unlabelled edges have a weight of $1$.
[…]
@@
Intuitively, it suffices to construct a strategy profile resulting in the play $v_{0}v_{1}v_{2}t_{1}v_{2}v_{1}v_{0}t_{2}^{\omega}$ such that the decisions of $\mathcal{P}_{1}$ in $v_{3}$ prevent $\mathcal{P}_{3}$ and $\mathcal{P}_{4}$ from having a profitable deviation. This can be ensured by having $\mathcal{P}_{1}$ avoiding the target of the owner of the vertex from which $v_{3}$ was reached (as they must have deviated). Formally, we consider a strategy $\sigma_{1}$ of $\mathcal{P}_{1}$ such that $\sigma_{1}(v_{0})=v_{1}$, $\sigma_{1}(v_{0}v_{1}v_{2}t_{1}v_{2}v_{1}v_{0})=t_{2}$ and for all histories $h$ such that $\mathsf{last}(h)=v_{3}$, we have $\sigma_{1}(h)=t_{4}$ if the penultimate vertex of $h$ is $v_{1}$ and $\sigma_{1}(h)=t_{3}$ otherwise (in particular, when the penultimate vertex of $h$ is $v_{2}$). We let $\sigma_{2}$ denote the unique strategy of $\mathcal{P}_{2}$ in the arena. For $\mathcal{P}_{3}$, we consider a strategy $\sigma_{3}$ such that $\sigma_{3}(v_{0}v_{1})=v_{2}$ and $\sigma_{3}(v_{0}v_{1}v_{2}t_{1}v_{2}v_{1})=v_{0}$. Finally, for $\mathcal{P}_{4}$, we let $\sigma_{4}$ be such that $\sigma_{4}(v_{0}v_{1}v_{2})=t_{1}$ and $\sigma_{4}(v_{0}v_{1}v_{2}t_{1}v_{2})=v_{1}$. The strategy profile $\sigma=(\sigma_{1},\sigma_{2},\sigma_{3},\sigma_{4})$ is an NE from $v_{0}$ with an outcome satisfying the objectives of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. The two players whose objectives are satisfied have no profitable deviations. The other two players can only visit their target by going through $v_{3}$ and the strategy of $\mathcal{P}_{1}$ prevents any deviation moving to $v_{3}$ from being profitable.
@@
The previous example shows that for Pareto optimal NEs (i.e., NEs whose cost profiles cannot be improved for a player without harming the cost of another), memory is necessary in general.
@@
In a zero-sum game, two players compete with opposing goals. Formally, a two-player zero-sum game is a two-player game $\mathcal{G}=(\mathcal{A},(\mathsf{cost}_{1},\mathsf{cost}_{2}))$ where $\mathcal{A}$ is a two-player arena and $\mathsf{cost}_{2}=-\mathsf{cost}_{1}$. We usually shorten the notation of a zero-sum game to $\mathcal{G}=(\mathcal{A},\mathsf{cost}_{1})$ due to the definition.
@@
[…]
Figure 2: An infinite weighted arena where there is no $\mathcal{P}_{2}$ optimal strategy from $v_{\infty}$ in the zero-sum shortest path game with $T=\{t\}$. All edges have a weight of $1$.

### v3

@@ Abstract
keywords:
cite=DBLP:conf/stacs/Main24
@@ 1 Introduction
Models in which players are not competing are also relevant for controller design: these are called non-zero-sum games, in which there can be more than two players. For instance, if the goal is to control a system consisting of several components, each with its own objective, it may prove too restrictive to assume that all components are adversarial to one another. In this work, we focus on Nash equilibria [29], a classical formalisation of rational behaviour in multiplayer games. Intuitively, a Nash equilibrium is a contract between the players, described by one strategy per player, such that no player can benefit by unilaterally breaking it. In the context of controller synthesis, as mentioned above, we are interested in finite-memory Nash equilibria. The goal of this work is to understand how much memory is needed for Nash equilibria in given classes of games.
@@
We study games on possibly infinite arenas in which all players share objectives of the same type. We focus on variants of reachability objectives; reachability objectives constitute a fundamental class of objectives in reactive synthesis [12]. A reachability objective models the goal of visiting a given set of target vertices. The complement of a reachability objective is called a safety objective: it requires that a set of undesirable vertices (e.g., unsafe configurations of the system) be avoided. A Büchi objective requires that a target set be visited infinitely often instead of only once as for reachability. The complement of a Büchi objective is called a co-Büchi objective; it models the requirement that undesirable configurations eventually no longer occur.
We also study games with a quantitative variant of the reachability objective. We consider arenas with edges labelled by non-negative integer weights (modelling, e.g., the passage of time), and define the cost of a player as the sum of weights prior to the first visit of a target, or infinity if the target is never reached. This cost function is called the shortest-path cost function. The goal of a player is to minimise it (i.e., to reach a target as soon as possible).
NEs are guaranteed to exist in games with these objectives or cost functions. For instance, see [10, 17] for reachability games and for shortest-path games in finite arenas and [32] for safety, Büchi and co-Büchi games. We also show in Appendix B that there always exist an NE in shortest-path games on infinite arenas.
In finite arenas, it has also been shown that finite-memory NEs are guaranteed to exist in the games we consider (e.g., [10, 33, 9]). However, existing constructions yield finite-memory NEs whose memory has a size that depends on the size of the arena. Therefore, these constructions do not directly generalise to infinite arenas. The main idea of these approaches is as follows. First, one shows that there exist plays resulting from NEs with a finite representation, typically, a lasso. This play is then encoded in a Mealy machine. If some player deviates from the play, the other players switch to a (finite-memory) punishing strategy to sabotage the deviating player; this enforces the stability of the equilibrium. This punishment mechanism is inspired by the proof of the folk theorem for NEs in repeated games [20, 30].
@@
Second, for reachability, shortest-path and safety games, we provide memory bounds that are independent of the size of the arena and depend only on the number of players. For Büchi and co-Büchi games, we show that similar arena-independent memory bounds cannot be obtained: we provide a family of two-player games where NEs with an outcome winning for the second player require a memory whose size is linear in the size of the (finite) arena (Proposition 7.3). We summarise our memory upper bounds for each considered type of game in Table 1.
@@
Reachability $n^{2}$ (Thm. 5.16)
Shortest-path $n^{2}+2n$ (Thm. 5.22)
@@
Büchi $|V|+n^{2}+n$ (Thm. 7.7) Finite-memory (Thm. 7.7, 7.12)
Co-Büchi $|V|+2n$ (Thm. 7.16) Finite-memory (Thm. 7.16)
@@
We build finite-memory NEs by using a relaxation of the classical punishment mechanism. We build NEs from plays that can be decomposed into segments (e.g., finite histories, simple plays or ultimately periodic plays) – see Lemmas 5.3, 6.4, 7.5, 7.10 and 7.14. Intuitively, we construct finite-memory strategies in which the players keep track of a segment of the decomposition (the one that they are currently attempting to construct) and the latest player to have moved. While this memory structure does not allow players to detect deviations within a segment of the decomposition, we impose conditions on decompositions such that no such deviation is profitable. The players switch to punishing strategies whenever there is a player who moves to a vertex outside of the current segment; the players retain enough information to detect such deviations.
Combined with the aforementioned NE existence results, our results imply that there always exist finite-memory NEs in the games we study. Memory is necessary in general to construct an NE if we require that its outcome satisfies the objectives of some players (see Example 2.3); our results provide upper bounds on the memory needed in these cases. Whether there always exists a memoryless NE in the games we consider is a closely related open question.
@@
We refer to the survey [13] for an extensive bibliography on games played on finite graphs, to [12] for a survey centred around reachability games and to [19] as a general reference on games on graphs. We discuss three research directions related to this work.
@@ 2 Preliminaries
Let $(V,E)$ be a directed graph where $V$ is a (possibly infinite) set of vertices and $E\subseteq V\times V$ is an edge relation. For any $v\in V$, we write $\mathsf{Succ}_{E}(v)=\{v^{\prime}\in V\mid(v,v^{\prime})\in E\}$ for the set of successor vertices of $v$. Let $n\in\mathbb{N}$. An $n$-player arena (on $(V,E)$) is a tuple $\mathcal{A}=((V_{i})_{i\in\llbracket{}n\rrbracket{}},E)$, where $(V_{i})_{i\in\llbracket{}n\rrbracket{}}$ is a partition of $V$ and, for all $i\in\llbracket{}n\rrbracket{}$, we say that the vertices in $V_{i}$ are controlled by $\mathcal{P}_{i}$. We assume that there are no deadlocks in the arenas we consider, i.e., for all $v\in V$, $\mathsf{Succ}_{E}(v)$ is not empty. We write $\mathcal{P}_{i}$ for player $i$.
@@
A play or history is simple if no vertex occurs twice (or more) within. A cycle is a history of the form $h=h^{\prime}\mathsf{first}(h^{\prime})$ where $h^{\prime}$ is a history, and $h=h^{\prime}\mathsf{first}(h^{\prime})$ is a simple cycle if $h^{\prime}$ is a simple history. A play $\pi$ is a lasso if there exist $p\in V^{*}$ and $c\in\mathsf{Hist}(\mathcal{A})$ such that $\pi=pc^{\omega}$. For any $p\in V^{*}$ and $c\in\mathsf{Hist}(\mathcal{A})$ such that $pc^{\omega}\in\mathsf{Plays}(\mathcal{A})$, we say that the lasso $pc^{\omega}$ is simple if $pc$ is a simple history.
A segment of a play $\pi$ is either a suffix $\pi{\ell}$ of $\pi$ for some $\ell\in\mathbb{N}$ or any history of the form $v_{\ell}\ldots v_{\ell^{\prime}}$ where $\ell\leq\ell^{\prime}$ (i.e., an infix of $\pi$). We denote segments by $\mathsf{sg}$ to avoid distinguishing finite and infinite segments of plays in the following. A segment is simple if it is a simple history, a simple play or a simple lasso. A segment is trivial if it consists of a single vertex.
@@
There are two classes of strategies of interest in this work. A strategy $\sigma_{i}$ is memoryless if the moves it prescribes depend only on the current vertex, i.e., if for all $h$, $h^{\prime}\in\mathsf{Hist}_{i}(\mathcal{A})$, if $\mathsf{last}(h)=\mathsf{last}(h^{\prime})$, then $\sigma_{i}(h)=\sigma_{i}(h^{\prime})$. We view memoryless strategies as functions $V_{i}\to V$.
@@
To describe the strategy induced by a Mealy machine, we first define the iterated update function $\widehat{\mathsf{up}}\colon V^{*}\to M$ by induction. We write $\varepsilon$ for the empty word. We let $\widehat{\mathsf{up}}(\varepsilon)=m_{\mathsf{init}}$ and for all $w\in V^{*}$ and $v\in V$, we let $\widehat{\mathsf{up}}(wv)=\mathsf{up}(\widehat{\mathsf{up}}(w),v)$. The strategy $\sigma_{i}^{\mathcal{M}}$ induced by $\mathcal{M}$ is defined, for all histories $h=h^{\prime}v\in\mathsf{Hist}_{i}(\mathcal{A})$, by $\sigma_{i}^{\mathcal{M}}(h)=\mathsf{nxt}_{i}(\widehat{\mathsf{up}}(h^{\prime}),v)$.
We say that a finite-memory strategy $\sigma_{i}$ has memory size at most $b\in\mathbb{N}$ if there is some Mealy machine $(M,m_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$ inducing $\sigma_{i}$ with $|M|\leq b$.
@@
In finite arenas, these two variants of Mealy machines share the same expressiveness. Given an edge-update Mealy machine, we can transform it into an equivalent vertex-update Mealy machine by augmenting the memory states with the last vertex to have been visited. With this last vertex and the current one, the last edge can be inferred. This allows memory updates to be imitated one step later than in the original edge-update Mealy machine. In general, edge-update Mealy machines may be more concise than vertex-update ones as they have access to more information in memory updates. It can be shown that transforming an edge-update Mealy machine into an equivalent vertex-update Mealy machine may require a vertex-update Mealy machine whose size is proportional to the product of the sizes of the edge-update Mealy machine size and the arena.
@@
We formalise the goal of a player in two ways. In the qualitative case, we describe the goal of a player by a set of plays, called an objective. We say that a play $\pi$ satisfies an objective $\Omega$ if $\pi\in\Omega$. For quantitative specifications, we assign a quantity to each play using a cost function $\mathsf{cost}_{i}\colon\mathsf{Plays}(\mathcal{A})\to\overline{\mathbb{R}}$ that $\mathcal{P}_{i}$ intends to minimise. Any goal expressed by an objective $\Omega$ can be encoded using a cost function $\mathsf{cost}_{i}$ which assigns 0 to plays in $\Omega$ and $1$ to others (i.e., the indicator of $\mathsf{Plays}(\mathcal{A})\setminus\Omega$); aiming to minimise this cost is equivalent to aiming to satisfy the objective. To avoid redundancy, we present definitions that are applicable for both objectives and cost functions by using cost functions, and separately present notions that are specific to objectives.
@@
We study five classes of games with goals related to reachability. Let $T\subseteq V$ denote a set of target vertices. We refer to the set of target vertices $T$ as a target.
The reachability objective formalises the goal of reaching the target. Formally, the reachability objective (for $T$) $\mathsf{Reach}(T)$ is defined by $\{v_{0}v_{1}v_{2}\ldots\in\mathsf{Plays}(\mathcal{A})\mid\exists\,\ell\in\mathbb{N},\,v_{\ell}\in T\}$. The complement of the reachability objective $\mathsf{Safe}(T)=\mathsf{Plays}(\mathcal{A})\setminus\mathsf{Reach}(T)$ is called the safety objective (for $T$) which requires avoiding $T$.
Next, we introduce a cost function for the goal of reaching a target as soon as possible. In this context, we assign (non-negative integer) weights to edges via a weight function $w\colon E\to\mathbb{N}$, which model, e.g., the time taken when traversing an edge. The weight function is extended to histories as follows: for $h=v_{0}\ldots v_{r}\in\mathsf{Hist}(\mathcal{A})$, we let $w(h)=\sum_{\ell=0}^{r-1}w((v_{\ell},v_{\ell+1}))$. We define the shortest-path cost function (for $T$ and $w$), for all plays $\pi=v_{0}v_{1}\ldots\in\mathsf{Plays}(\mathcal{A})$, by $\mathsf{SPath}^{T}_{w}(\pi)=w(\pi{r})$ if $r=\min\{\ell\in\mathbb{N}\mid v_{\ell}\in T\}$ exists and $\mathsf{SPath}^{T}_{w}(\pi)=+\infty$ otherwise. By assigning an infinite cost to plays that do not visit $T$, we model the idea that not reaching $T$ is the least desirable situation.
Finally, the Büchi objective requires reaching the target infinitely often. Formally, the Büchi objective (for $T$) $\mathsf{B\ddot{u}chi}(T)$ is defined by $\{v_{0}v_{1}v_{2}\ldots\in\mathsf{Plays}(\mathcal{A})\mid\forall\,\ell\in\mathbb{N},\,\exists\,\ell^{\prime}\geq\ell,\,v_{\ell^{\prime}}\in T\}$. The complement of a Büchi objective is a co-Büchi objective: the co-Büchi objective (for $T$) is defined as $\mathsf{coB\ddot{u}chi}(T)=\mathsf{Plays}(\mathcal{A})\setminus\mathsf{B\ddot{u}chi}(T)$ and requires visiting $T$ finitely often.
@@
NEs are guaranteed to exist in the five classes of games we consider; see, e.g. [10, 17] for reachability games and for shortest-path games in finite arenas, Appendix B for shortest-path games in general and [32] for safety, Büchi and co-Büchi games. Even though NEs exist from all initial vertices, there are no guarantees on the quality of their cost profiles (this is a core motivation of the constrained NE existence problem mentioned in Section 1). In a given game, there can exist several NEs from a given initial vertex with incomparable cost profiles. Similarly, in a given game with objectives, from a fixed initial vertex, there can exist NEs with which all players satisfy their objective and NEs with which no players satisfy their objective [32, 33]. In general, memory may be necessary to construct an NE with a cost profile that is preferable to a given vector (i.e., a solution to the constrained existence problem).
@@
$v_{0}$$t_{12}$$v_{1}$$v_{2}$$t_{1}$33
(a) A two-player weighted arena in which there are two memoryless NE with incomparable payoff profiles. Edges are labelled by their weight and unlabelled edges have a weight of $1$.
$v_{0}$$t_{2}$$v_{1}$$v_{2}$$t_{1}$$v_{3}$$t_{3}$$t_{4}$
@@
Intuitively, it suffices to construct a strategy profile resulting in the outcome $v_{0}v_{1}v_{2}t_{1}v_{2}v_{1}v_{0}t_{2}^{\omega}$ such that the decisions of $\mathcal{P}_{1}$ in $v_{3}$ prevent $\mathcal{P}_{3}$ and $\mathcal{P}_{4}$ from having a profitable deviation. This can be ensured by having $\mathcal{P}_{1}$ avoiding the target of the owner of the vertex from which $v_{3}$ was reached (as they must have deviated). Formally, we consider a strategy $\sigma_{1}$ of $\mathcal{P}_{1}$ such that $\sigma_{1}(v_{0})=v_{1}$, $\sigma_{1}(v_{0}v_{1}v_{2}t_{1}v_{2}v_{1}v_{0})=t_{2}$ and for all histories $h$ such that $\mathsf{last}(h)=v_{3}$, we have $\sigma_{1}(h)=t_{4}$ if the penultimate vertex of $h$ is $v_{1}$ and $\sigma_{1}(h)=t_{3}$ otherwise (in particular, when the penultimate vertex of $h$ is $v_{2}$). We let $\sigma_{2}$ denote the unique strategy of $\mathcal{P}_{2}$ in the arena. For $\mathcal{P}_{3}$, we consider a strategy $\sigma_{3}$ such that $\sigma_{3}(v_{0}v_{1})=v_{2}$ and $\sigma_{3}(v_{0}v_{1}v_{2}t_{1}v_{2}v_{1})=v_{0}$. Finally, for $\mathcal{P}_{4}$, we let $\sigma_{4}$ be such that $\sigma_{4}(v_{0}v_{1}v_{2})=t_{1}$ and $\sigma_{4}(v_{0}v_{1}v_{2}t_{1}v_{2})=v_{1}$. The strategy profile $\sigma=(\sigma_{1},\sigma_{2},\sigma_{3},\sigma_{4})$ is an NE from $v_{0}$ with an outcome satisfying the objectives of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. The two players whose objectives are satisfied have no profitable deviations. The other two players can only visit their target by going through $v_{3}$ and the strategy of $\mathcal{P}_{1}$ prevents any deviation moving to $v_{3}$ from being profitable.
@@
The previous example shows that for Pareto optimal NEs (i.e., NEs whose cost profiles cannot be improved for a player without harming the cost of another), memory is necessary in general. Whether there always exists a memoryless NE in the games we consider is an open question.
@@
In a zero-sum game, two players compete with opposing goals. Formally, a two-player game $\mathcal{G}=(\mathcal{A},(\mathsf{cost}_{1},\mathsf{cost}_{2}))$ is a zero-sum game if $\mathsf{cost}_{2}=-\mathsf{cost}_{1}$. We usually shorten the notation of a zero-sum game to $\mathcal{G}=(\mathcal{A},\mathsf{cost}_{1})$ due to the definition.
@@
$_{0}vv_{\infty}$$v_{1}$$v_{2}$$v_{3}$$t$
Figure 2: An infinite weighted arena where there is no $\mathcal{P}_{2}$ optimal strategy from $v_{\infty}$ in the zero-sum shortest path game with $T=\{t\}$. All edges have a weight of $1$. Circle and squares respectively denote $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ vertices.